OpenStudy (anonymous):

3 * 6x+7y - 2(sqrt)7= (6x-2y)+4 can someone answer this?

OpenStudy (anonymous):

3(6x)+7y-2/7=(6x-2y)+4 distribute 18x+7y-2/7=6x-2y+4 move 18x to the other side 7y-2/7=-12x-2y+4 add 2/7 7y=-12x-2y+4+2/7 add 2y 9y=-12x+4 2/7 divide by 9 y=-4/3x+10/21 solved for y 18x+7y-2/7=6x-2y+4 subtract 7y 18x-2/7=6x-5y+4 add 2/7 18x=6x-5y+(30/7) subtract 6x 12x=-5y+(30/7) divide by 12 x=-5/12y+5/14 solved for x